40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 24 11:31:59 2018, gautam, Configuration, ALS, Proposed changes to EX fiber coupling EX_fiber_MM.pdfEX_fiber_changes.png
    Reply  Tue Apr 24 18:54:15 2018, gautam, Configuration, ALS, Proposed changes to EX fiber coupling 
       Reply  Wed Apr 25 19:09:37 2018, gautam, Configuration, ALS, New look EX Fiber coupling EX_fiber_MM.pdfIMG_6977.JPGIMG_6972.JPGIMG_6974.JPGIMG_6976.JPG
          Reply  Thu Apr 26 11:24:50 2018, gautam, Configuration, ALS, New look EX Fiber coupling - pol stability polStab.pdf
             Reply  Thu Apr 26 18:58:21 2018, Bruce, Configuration, ALS, New look EX Fiber coupling - pol stability 
             Reply  Fri Apr 27 01:36:02 2018, gautam, Configuration, ALS, IR ALS noise performance 2018_04_BeatMouth_POX.pdf2018_04_BeatMouth.pdfALSSpecgram.pdfALS_ASD.pdfALSstab.pdf
                Reply  Wed May 2 21:39:33 2018, gautam, Configuration, ALS, IR ALS for EY 
                   Reply  Fri May 4 21:17:57 2018, gautam, Configuration, ALS, BeathMouth pulled out of PSL table 
                      Reply  Tue May 8 00:40:51 2018, gautam, Configuration, ALS, BeathMouth pulled out of PSL table IMG_6986.JPGIMG_6987.JPG
                         Reply  Thu May 24 19:41:28 2018, gautam, Configuration, ALS, BeathMouth reinstalled on PSL table 
                      Reply  Thu May 24 13:06:17 2018, gautam, Configuration, ALS, DFD noises DFDcalib.pdfDFD_NB.pdf
                         Reply  Thu May 24 20:31:03 2018, gautam, Configuration, ALS, DFD noises TF_X_mag.pdfTF_X_phase.pdf
Message ID: 13796     Entry time: Fri Apr 27 01:36:02 2018     In reply to: 13791     Reply to this: 13807
Author: gautam 
Type: Configuration 
Category: ALS 
Subject: IR ALS noise performance 

Summary:

My goal was to do some further characterization of the IR ALS system tonight. With POX as an OOL sensor, I measured an RMS displacement noise of  8 pm with the arm under ALS control. I calculated the CARM linewidth to be 77 Hz (=10.3 pm) for the 40m parameters, assuming 30ppm arm loss. Fuurthermore, this number is 3x better than the 24 pm RMS quoted in the Izumi et. al. paper. Of course I am quoting the best results from my efforts tonight. Conclusions:

  1. [Attachment #1] --- With XARM locked using POX, the ALS beat noise (i.e. Phase Tracker output noise) lines up well with the reference we have been using for some time now (and indeed, is better in some places).
  2. [Attachment #2] --- With the arm locked on ALS and POX as an OOL sensor, I measured performance comparable to this measurement we did sometime last year. Anomalies in this measurement and the one above were what precipitated the IMC noise investigation.
  3. [Attachment #3] --- The above two attachments are not the whole story. During the day, I get significantly worse performance (so much so that I couldn't even do the handoff to ALS control). But in 5 minutes of measurement, the ALS noise seems quite stationary.
  4. [Attachment #4] --- This is really the same as Attachment #2, but I wanted to overlay some vlines. Maybe this is a clue to some 60 Hz / ground loop issues, but the RMS has significant contribution from these harmonics. Tmrw, I will add the old measurement overlaid to this plot (and for what its worth, the Izumi et. al. spectrum as well).
  5. [Attachment #5] --- With the arm under ALS control, I was able to maintain the lock for a solid hour (and more as I write up this elog). Somehow inkscape screwed up the fonts, but main point here is that TRX is stable to within 10% throughout the observation time.

Since the stability and noise seemed quite good, I decided to collect some arm scan data to give to our modeSpec SURFs to practice fitting (which is the short dip in TRX in Attachment #4). Although after the discussion with Rana today, I think it may be that we want to do this measurement in reflection and not transmission, and look for a zero crossing in the PDH signal. In any case, I was able to scan 7 FSRs without any issues. I will upload the data to some git repo. GPS start time is 1208850775, sweep was 3mins long.

I think the next step here is to noise-budget this curve. At least the DFD noises

Attachment 1: 2018_04_BeatMouth_POX.pdf  39 kB  Uploaded Fri Apr 27 02:49:25 2018  | Hide | Hide all | Show all
2018_04_BeatMouth_POX.pdf
Attachment 2: 2018_04_BeatMouth.pdf  17 kB  Uploaded Fri Apr 27 02:49:34 2018  | Hide | Hide all | Show all
2018_04_BeatMouth.pdf
Attachment 3: ALSSpecgram.pdf  265 kB  Uploaded Fri Apr 27 02:58:28 2018  | Show | Hide all | Show all
Attachment 4: ALS_ASD.pdf  234 kB  Uploaded Fri Apr 27 02:58:38 2018  | Show | Hide all | Show all
Attachment 5: ALSstab.pdf  78 kB  Uploaded Fri Apr 27 02:58:51 2018  | Hide | Hide all | Show all
ALSstab.pdf
ELOG V3.1.3-